Puleng receives R1500,00 from a bank now, that charges 10.5% simple discount per annum. She has to pay back an amount of R1893.75 in a few months time. The number of months after which Puleng pays back the loan, rounded to one decimal place is

(1) 2.5

(2) 23.8

(3) 2.0

(4) 30.0


1
Expert's answer
2022-02-15T11:05:24-0500

Solution:

"P=R\ 1500
\\r=10.5\%=0.105
\\A=R\ 1893.75"

No. of months = t = ?

"A=P+\dfrac{Prt}{12}
\\\Rightarrow 1893.75=1500+\dfrac{1500\times0.105\times t}{12}
\\\Rightarrow 393.75=\dfrac{1500\times0.105\times t}{12}
\\ \Rightarrow t=30"

So, option 4 is correct.